Bianca lives in Arizona and earned $2350$2350 in taxable income this biweekly pay period. The table shows tax withholdings for a biweekly pay period. Note that in Arizona, individuals elect to withhold a certain percentage of their taxable income; the withholding does not involve tax brackets.
Tax Amount
Social Security 6.2%6.2% of taxable income
Medicare 1.45%1.45% of taxable income
Federal income tax withholding $174.70$174.70 plus 22%22% over $1664$1664
State income tax withholding 3.6%3.6% of taxable income
Using the table, what is the amount withheld from Bianca's paycheck? Enter your answer as a number rounded to the nearest dollar, such as: $425.
Answer #1
Social Security withheld 146 =2350*6.2%
Medicare withheld 34 =2350*1.45%
Federal income tax withholding 326 =174.70+(2350-1664)*22%
State income tax withholding 85 =2350*3.6%
Amount withheld from Bianca's paycheck 590 or 591